Determine the pressure, in psia, 60 ft below the surface of the sea. (The specific gravity of the sea is 1.022)
1
Expert's answer
2021-10-31T18:11:45-0400

The pressure is given as follows:


"p = \rho g h"

where "\rho = 1022kg/m^3" is the specific gravity of the seawater, "g = 9.81m\/s^2" is the gravitational acceleration, and "h = 60ft\approx 18.29m" is the depth. Thus, obtain:


"p = 1022\cdot 9.81\cdot 18.29 \approx 1.834\times 10^5Pa"

In psia:


"p_a = p/6895 + 14.7psi\approx 41.29psia"

where 14.7psi is the atmospheric pressure.


Answer. 41.29 psia.